What is the [CH3CO2-]/[CH3CO2H] ratio necessary to make a buffer solution with a pH of 4.44? The Ka for acetic acid is 1.76 x 10^-5.

The answer choices are

A) .39:1

B) .49:1

C) 1.07:1

D) 1.33:1

E) 4.44:1

It is NOT A according to my professor counting that wrong on the exam, I just want the answer and how it works please!
Answer #1

Solution:

pH of an acidic buffer is calculated by Henderson equation as:

pH = pKa + log [ salt] / [acid]

Given, ka for acetic acid = 1.76 x 10^-5

Therefore, pKa = -log ka = - log 1.76 x 10^-5

= 5 - log 1.76 = 5 - 0.246 = 4.75

pH = 4.44

Hence, using Henderson equaion,

pH = pka + log [CH3COO- ] /[CH3COOH]

4.44 = 4.75 + log [CH3COO- / CH3COOH]

log [CH3COO- / CH3COOH ] = - 0.31

[CH3COO- / CH3COOH] = antilog -0.31 = 0.49

Therefore, ratio of ,

[CH3COO- / CH3COOH] = 0.49 : 1

Hence, option B.